# TAPTRv3: Spatial and Temporal Context Foster Robust Tracking of Any Point in Long Video

### v1

@@ 3.2 Visibility-aware Long-Temporal Attention
The RNN-like long-temporal modeling in TAPTR and TAPTRv2 can lead to feature drift, primarily due to the excessive feature updates during testing, rooted in the disparity in video lengths between training (24 frames) and testing (ranging from 50 to 1300 frames). Thus, as shown in Fig. 3 (b), we resort to the attention mechanism, for its ability to handle varying token lengths, as demonstrated in modern LLMs [51, 52, 31, 10]. Following modern LLMs, we also utilize the rotary positional embedding [37] to help our long-temporal attention pay more attention to recent frames. More specifically, the long-temporal attention weights of the target tracking point between the $t$-th frame and all past frames can be first computed as,
$\begin{gathered}\mathbf{F}_{t}=\left[\mathbf{f}_{0},\mathbf{f}_{1},\cdots,\mathbf{f}_{t-1}\right]^{\top},\mathbf{R}_{t}=\left[\mathbf{r}_{0},\mathbf{r}_{1},\dots,\mathbf{r}_{t-1}\right]^{\top},\\ \mathbf{d}_{t}^{\prime}=\texttt{SoftMax}\left(\left(\mathbf{F}_{t}+\mathbf{R}_{t}\right)\otimes\left(\mathbf{f}_{t}^{\prime}+\mathbf{r}_{t}\right)\right),\\ \end{gathered}$ (2)
@@
where $\mathbf{F}_{t}\in\mathbb{R}^{t\times D}$ and $\mathbf{R}_{t}\in\mathbb{R}^{t\times D}$ are the point query’s refined content features in the past frames of the $t$-th frame, and their corresponding rotary positional embeddings, $\mathbf{d}_{t}^{\prime}\in\mathbb{R}^{t}$ is the long-temporal attention distribution (weights), $\otimes$ indicates the matrix multiplication, and $\mathbf{f}_{t}^{\prime}\in\mathbb{R}^{D}$ is the content feature of the point query in the $t$-th frame that has not been fully refined by decoder. In the first decoder layer, $\mathbf{f}_{t}^{\prime}\equiv\mathbf{f}$.
@@
Different from the textual scenarios, since the target tracking point will be occluded sometimes, the refined content features from the frames that the target tracking point is occluded may contribute noise to the long-temporal attention. To prevent being affected by the noise, we utilize the visibility predictions in the past frames $\mathbf{a}_{t}\in\mathbb{R}^{t}=\left[\alpha_{0},\alpha_{1},\dots,\alpha_{t-1}\right]^{T}$ to reweight the attention distribution, making it visibility-aware. The final visibility-aware attention weights are used to weight-sum their corresponding content features to obtain temporal querying result $\Delta\mathbf{f}_{t}^{\mathcal{T}}\in\mathbb{R}^{D}$. The querying result will be further utilized as a residual to update the point query’s content feature to complete the VLTA. The process can be formulated as,
@@
$\begin{gathered}\mathbf{d}_{t}=\frac{\mathbf{d}_{t}^{\prime}\odot\mathbf{a}_{t}}{\texttt{Sum}(\mathbf{a}_{t})},\Delta\mathbf{f}_{t}^{\mathcal{T}}=\mathbf{F}_{t}^{\top}\otimes\mathbf{d}_{t},\\ \mathbf{f}_{t}^{\prime}\Leftarrow\texttt{LN}\left(\mathbf{f}_{t}^{\prime}+\Delta\mathbf{f}_{t}^{\mathcal{T}}\right),\end{gathered}$ (3)
@@
where $\mathbf{d}_{t}\in\mathbb{R}^{t}$ is the attention distribution that is reweighted by visibility predictions, $\mathbf{f}_{t}^{\prime}$ is the output and will be sent to the following modules for further refinement, LN is the Layernorm [21], and $\odot$ is the element-wise multiplication.
@@ 3.3 Context-aware Cross-Attention with APU
Unlike object-level DETR-like methods, the content feature of a point query in TAPTRv3 is a point-level feature. This granularity can limit the model’s receptive field during cross-attention with the image feature map, often causing ambiguity in attention weights. This issue becomes more serious when the target tracking point undergoes significant varia-
@@
tions or when the image contains uniform regions or repetitive patterns. Such conditions can lead to noisy querying of spatial features as well as noisy position updates in the cross-attention’s belonging APU block. Inspired by previous optical flow estimation methods [39] and the recent LocoTrack [5], we propose to leverage more informative spatial context features to provide the point query with a more comprehensive understanding of its surrounding environment, thus helping eliminate ambiguity.
As illustrated in Fig. 3 (c), different from the vanilla key-aware deformable attention [23], the query’s point-level content feature will only be used to predict $M$ sampling offsets $\mathbf{O}_{t}\in\mathbb{R}^{M\times 2}=\left[\mathbf{o}_{t}^{0},\mathbf{o}_{t}^{1},\dots,\mathbf{o}_{t}^{M-1}\right]^{\top}$ by an MLP. For the corresponding attention weights, we obtain them by leveraging the patch-level context features. More specifically, as shown in Fig. 4, for the $m$-th sampling point, its corresponding spatial context features in the $t$-th frame $\mathbf{K}_{t}^{m}\in\mathbb{R}^{N^{2}\times D}$ can be constructed by
$\begin{gathered}\mathbf{K}_{t}^{m}=\texttt{Bili}\left(\mathbf{X}_{t},\mathbf{l}_{t}^{\prime}+\mathbf{o}_{t}^{m}+\mathbf{G}\right),\\ \end{gathered}$ (4)
where $\mathbf{l}_{t}^{\prime}\in\mathbb{R}^{2}$ is the positional part of the point query in the $t$-th frame that has not been fully refined by decoder layers. In the first decoder layer’s CCA module $\mathbf{l}_{t}^{\prime}\equiv\mathbf{l}_{t-1}$. After that, the $m$-th sampling point’s corresponding patch-level similarity $w_{t}^{m}\in\mathbb{R}$ is calculated as,
$\begin{gathered}\mathbf{S}_{t}^{m}\in\mathbb{R}^{N^{2}\times N^{2}}=\mathbf{C}\otimes{\mathbf{K}_{t}^{m}}^{\top},\\ w_{t}^{m}=\texttt{MLP}(\texttt{Flatten}(\mathbf{S}_{t}^{m})),\end{gathered}$ (5)
where $\mathbf{S}_{t}^{m}$ is the intermediate representation of the patch-level similarity, and Flatten indicates the flatten operation. By sending the intermediate representation $\mathbf{S}_{t}^{m}$ to an MLP, the more fine-grained similarities between every two points in the two patches are comprehensively considered, leading to a high-quality patch-level similarity. The patch-level similarities between the point query and all sampling points $\mathbf{w}_{t}\in\mathbb{R}^{M}=[w_{t}^{0},w_{t}^{1},\dots,w_{t}^{M-1}]^{\top}$ will work as attention weights to aggregate their corresponding values $\mathbf{V}_{t}\in\mathbb{R}^{M\times D}$, obtaining the spatial querying result $\Delta\mathbf{f}_{t}^{\mathcal{S}}\in\mathbb{R}^{D}$,
$\begin{gathered}\mathbf{V}_{t}=[\mathbf{v}_{t}^{0},\mathbf{v}_{t}^{1},\cdots,\mathbf{v}_{t}^{M-1}]^{\top},\mathbf{v}_{t}^{m}=\texttt{Bili}(\mathbf{X}_{t},\mathbf{l}_{t}^{\prime}+\mathbf{o}_{t}^{m}),\\ \Delta\mathbf{f}_{t}^{\mathcal{S}}=\mathbf{V}_{t}^{\top}\otimes\texttt{SoftMax}\left(\mathbf{w}_{t}/\sqrt{D}\right),\\ \end{gathered}$ (6)
where $\mathbf{V_{t}}$ are the sampled values of $M$ sampling points on $\mathbf{X}_{t}$. After that, the spatial querying results will be utilized as a residual to complete the CCA,
@@
Following TAPTRv2, we also add the attention-based position update (APU) in CCA, to benefit from the attention weights’ resilience to the domain gap without contaminating the point query. This process can be formulated as,
$\begin{gathered}\Delta\mathbf{l}_{t}=\mathbf{O}_{t}^{\top}\otimes\texttt{SoftMax}\left(\texttt{MLP}\left(\mathbf{w}_{t}\right)/\sqrt{D}\right),\\ \mathbf{l}_{t}^{\prime}\Leftarrow\mathbf{l}_{t}^{\prime}+\Delta\mathbf{l}_{t},\end{gathered}$ (8)
@@
3.4 Tracking Reestablish with Global Matching
In general, point motions are smooth in video. However, in long videos, especially in offline scenarios, scene cut often occurs, leading to sudden large motions. Since the initial position of a point query in the current frame is inherited from the previous frame’s prediction as we have described in Sec. 3.1, the sudden large motion will cost many frames for TAPTRv3 to catch up with the target tracking point.
Thus, when a scene cut occurs [^5] We use the off-the-shelf PySceneDetect to detect the scene cuts. at the $t$-th frame, we trigger the global matching module to help TAPTRv3 reestablish tracking. Similar to previous methods [5, 7], the global matching module will construct a similarity map $\mathbf{H}_{t}\in\mathbb{R}^{H\times W}$ between the target tracking point and the current frame’s feature map. However, instead of relying on a point-level feature as in previous methods, similar to CCA, we leverage the spatial context features to help improve the accuracy of the similarity map. After that, SoftArgMax is conducted on the similarity map to obtain a final position prediction $\mathbf{l}_{t}\in\mathbb{R}^{2}$ that is differentiable. The prediction will be used to replace the unreliable position prediction of the current frame from the transformer decoder. The replacement helps TAPTRv3 to prevent wrong predictions in the subsequent frames. The process can be formulated as,
@@ 4 Experiments
Training Kinetics RGB-Stacking RoboTAP DAVIS
Method Data AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA AJ $<\delta^{x}_{avg}$ OA
PIPs [12] FT 31.7 53.7 72.9 59.1 42.2 64.8 77.7
PIPs++ [54] PO 63.5 58.5 63.0 69.1
TAP-Net [6] Kub24 38.5 54.4 80.6 53.5 68.1 86.3 45.1 62.1 82.9 33.0 48.6 78.8
TAPIR [7] Kub24 49.6 64.2 85.0 55.5 69.7 88.0 59.6 73.4 87.0 56.2 70.0 86.5
CoTracker [20] Kub24 49.6 64.3 83.3 67.4 78.9 85.2 58.6 70.6 87.0 61.8 76.1 88.3
TAPTR [26] Kub24 49.0 64.4 85.2 60.8 76.2 87.0 60.1 75.3 86.9 63.0 76.1 91.1
TAPTRv2 [25] Kub24 49.7 64.2 85.7 53.4 70.5 81.2 60.9 74.6 87.7 63.5 75.9 91.4
LocoTrack [5] Kub24 52.9 66.8 85.3 69.7 83.2 89.5 62.3 76.2 87.1 63.0 75.3 87.2
BootsTAPIR [8] Kub24+15M 54.6 68.4 86.5 70.8 83.0 89.9 64.9 80.1 86.3 61.4 73.6 88.7
CoTracker3 (online) [19] Kub64 54.1 66.6 87.1 71.1 81.9 90.3 60.8 73.7 87.1 64.5 76.7 89.7
CoTracker3 (online) [19] Kub64+15K 55.8 68.5 88.3 71.7 83.6 91.1 66.4 78.8 90.8 63.8 76.3 90.2
TAPTRv3 (Ours) Kub24 54.5 67.5 88.2 73.0 86.2 90.0 64.6 77.2 90.1 63.2 76.7 91.0
Table 1: Comparison of TAPTRv3 with prior methods. Note that, CoTracker3 is a concurrent work. CoTracker3 and BootsTAPIR both introduced extra data for training. Specifically, CoTracker3 re-rendered synthetic Kubric videos with length of 64 frames and additionally incorporated 15K real videos. In contrast, BootsTAPIR trained on an additional 15M real videos. TAPTRv3 obtains state-of-the-art performance on most datasets and remains competitive with methods trained on extra internal data. Training data: (Kub24) Kubric [11] with 24 frames per video, (Kub64) Kubric with 64 frames per video, (PO) PointOdyssey [54], (FT) FlyingThings++ [29].
We conduct extensive experiments on multiple challenging evaluation datasets collected from both the real and synthetic world to verify the superiority of TAPTRv3.
@@ 4.1 Datasets and Evaluation
Training Data. For fair comparison, following the previous methods [7, 20, 26, 25, 5], we trained our model on the TAP-Vid-Kubric [6] dataset, which consists of 11,000 synthetic videos generated by Kubric Engine [11], each containing 24 frames showing 3D rigid objects falling to the ground and bouncing. Each video has 2,048 points sampled on moving objects and backgrounds to be tracked, and their corresponding trajectories are also generated for training. The points to be tracked in the video are occasionally occluded to allow the model to cope with this situation. During training, the resolution of the videos is resized to $512\times 512$ and we randomly select 800 trajectories in each video for efficient training.
@@
Evaluation Data. We follow previous methods to evaluate TAPTRv3 on the challenging TAP-Vid [6] benchmark, which consists of 3 subsets, TAP-Vid-Kinetics, TAP-Vid-DAVIS, and RGB-Stacking. TAP-Vid-DAVIS comprises 30 real-world videos from the DAVIS 2017 validation set [32]. These videos are relatively short, averaging less than 70 frames. TAP-Vid-Kinetics contains 1,144 YouTube videos from the Kinetics-700-2020 validations set [4], with camera shakes, complex environments, and even scene cuts. These videos are relatively long, averaging about 250 frames per video. RGB-Stacking is a synthetic dataset that captures the process of a robotic arm grasping solid-colored blocks, with an average duration of about 250 frames. Although it has a relatively smaller domain gap compared to the training set, which is also a synthetic dataset, the objects in this dataset usually lack texture, making it also difficult to track. In addition, we also evaluated TAPTRv3 on RoboTAP [43], which comprises 265 real-world videos from robotic manipulation tasks. The video length in this dataset varies significantly, with the longest videos reaching up to 1300 frames and the average length of about 270 frames per video.
@@
Evaluation Metrics and Settings. We use the standard metrics proposed in TAP-Vid [6] for evaluation, including three important metrics. Occlusion Accuracy (OA), describes the accuracy of classifying whether a target tracking point is visible or occluded. $<\delta_{avg}^{x}$, reflecting the average precision of visible points’ position at thresholds of 1, 2, 4, 8, and 16 pixels. Average Jaccard (AJ), a comprehensive metric, considers both the precision of position and visibility prediction. Since TAPTRv3 is an online tracker, we use the “First query” mode [6] to evaluate the model, which tracks the target tracking point from the first frame when they are visible until the end of the video. This is much more difficult than the “Strided query” mode for offline trackers. Besides, since the resolution of the input video has a large impact on the final performance, we limit the resolution of the input video to $256\times 256$ for a fair comparison with other methods during evaluation.
Row LTA Vis. Aware No Window No Inv. Sup. CCA AJ $<\delta^{x}_{avg}$ OA
1 46.0 62.8 83.8
2 48.9 63.0 85.7
3 50.1 63.1 85.9
4 51.1 63.3 85.8
5 51.9 64.8 86.9
6 52.9 65.8 87.2
Table 2: Ablations on key component. “LTA” refers to Long-Temporal Attention, “Vis.” is short for visibility, “No Window” indicates the removal of the sliding window, and “No Inv. Sup.” represents no supervision of the invisible points’ positions.
Patch-level Similarity AJ $<\delta^{x}_{avg}$ OA
Element-wise 52.5 65.1 85.7
Every two point 52.9 65.8 87.2
Table 3: Ablation on patch-level similarity.
Update Methods AJ $<\delta^{x}_{avg}$ OA
VLTA 52.6 65.3 86.8
MLP 52.5 65.7 87.3
No Updates 52.9 65.8 87.2
Table 4: Ablation on context features updating.

### v2

@@ 3.2 Visibility-aware Long-Temporal Attention
The RNN-like long-temporal modeling in TAPTR and TAPTRv2 can lead to feature drift, primarily due to the excessive feature updates during testing, rooted in the disparity in video lengths between training (24 frames) and testing (ranging from 50 to 1300 frames). Thus, as shown in Fig. 1 (b), we resort to the attention mechanism for its ability to handle varying token lengths, as demonstrated in modern LLMs (Zhang et al., 2025; 2023b; OpenAI, 2023; Google, 2023). Following modern LLMs, we also utilize the rotary positional embedding (Su et al., 2024) to help our long-temporal attention pay more attention to recent frames. More specifically, the long-temporal attention weights of the target tracking point between the $t$-th frame and all past frames [^2] In practice, it is not necessary to interact with all past frames, and this is only for ease of description. For more details, please refer to the Sec. A.2 in our appendix. can be first computed as:
[…]
@@
where $\mathbf{F}_{t}\in\mathbb{R}^{t\times D}$ and $\mathbf{R}_{t}\in\mathbb{R}^{t\times D}$ are the point query’s refined content features in the past frames of the $t$-th frame, and their corresponding rotary frame index embeddings, $\mathbf{d}_{t}^{\prime}\in\mathbb{R}^{t}$ is the long-temporal attention distribution (weights), $\otimes$ indicates the matrix multiplication, and $\mathbf{f}_{t}^{\prime}\in\mathbb{R}^{D}$ is the content feature of the point query in the $t$-th frame that has not been fully refined by decoder. In the first decoder layer, $\mathbf{f}_{t}^{\prime}\equiv\mathbf{f}$.
@@
Different from the textual scenarios, since the target tracking point will be occluded sometimes, the refined content features from the frames in which the target tracking point is occluded may contribute noise to the long-temporal attention. To prevent being affected by the noise, we utilize the visibility predictions in the past frames $\mathbf{a}_{t}\in\mathbb{R}^{t}=\left[\alpha_{0},\alpha_{1},\dots,\alpha_{t-1}\right]^{T}$ to reweight the attention distribution, making it visibility-aware. The final visibility-aware attention weights are used to weight-sum their corresponding content features to obtain the temporal querying result $\Delta\mathbf{f}_{t}^{\mathcal{T}}\in\mathbb{R}^{D}$. The querying result will be further utilized as a residual to update the point query’s content feature to complete the VLTA. The process can be formulated as:
@@
$\begin{gathered}\mathbf{d}_{t}=\frac{\mathbf{d}_{t}^{\prime}\odot\mathbf{a}_{t}}{\texttt{Sum}(\mathbf{a}_{t})},\penalty 10000\ \penalty 10000\ \penalty 10000\ \Delta\mathbf{f}_{t}^{\mathcal{T}}=\mathbf{F}_{t}^{\top}\otimes\mathbf{d}_{t},\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbf{f}_{t}^{\prime}\Leftarrow\texttt{LN}\left(\mathbf{f}_{t}^{\prime}+\Delta\mathbf{f}_{t}^{\mathcal{T}}\right),\end{gathered}$ (3)
@@
where $\mathbf{d}_{t}\in\mathbb{R}^{t}$ is the attention distribution that is reweighted by visibility predictions, $\mathbf{f}_{t}^{\prime}$ is the output and will be sent to the following modules for further refinement, LN is the Layernorm (Lei Ba et al., 2016), and $\odot$ is the element-wise multiplication.
@@ 3.3 Context-aware Cross-Attention with APU
Unlike object-level DETR-like methods, the content feature of a point query in TAPTRv2 is a point-level feature. This granularity limits the model’s receptive field during cross-attention with the image feature map, often causing ambiguity in attention weights. This issue becomes more serious when the target tracking point undergoes significant variations or when the image contains uniform regions or repetitive patterns. Such conditions can lead to noisy querying of spatial features as well as noisy position updates in the cross-attention’s belonging APU block. Inspired by previous methods (Teed & Deng, 2020; Bian et al., 2023; Cho et al., 2024), we propose integrating richer spatial context into the attention mechanism. This provides point queries with a more comprehensive understanding of their surroundings, resulting in more accurate and robust attention weights.
@@
As illustrated in Fig. 1 (c), different from the vanilla key-aware deformable attention (Li et al., 2023), the query’s point-level content feature will only be used to predict $M$ sampling offsets $\mathbf{O}_{t}\in\mathbb{R}^{M\times 2}=\left[\mathbf{o}_{t}^{0},\mathbf{o}_{t}^{1},\dots,\mathbf{o}_{t}^{M-1}\right]^{\top}$ by an MLP. For the corresponding attention weights, we obtain them by leveraging the patch-level context features. More specifically, as shown in Fig. 2, for the $m$-th sampling point, its corresponding spatial context features in the $t$-th frame $\mathbf{K}_{t}^{m}\in\mathbb{R}^{N^{2}\times D}$ can be constructed by:
[…]
where $\mathbf{l}_{t}^{\prime}\in\mathbb{R}^{2}$ is the positional part of the point query in the $t$-th frame that has not been fully refined by decoder layers. In the first decoder layer’s CCA module $\mathbf{l}_{t}^{\prime}\equiv\mathbf{l}_{t-1}$. After that, the $m$-th sampling point’s corresponding patch-level similarity $w_{t}^{m}\in\mathbb{R}$ is calculated as:
$\begin{gathered}\mathbf{S}_{t}^{m}\in\mathbb{R}^{N^{2}\times N^{2}}=\mathbf{C}\otimes{\mathbf{K}_{t}^{m}}^{\top},\penalty 10000\ \penalty 10000\ \penalty 10000\ w_{t}^{m}=\texttt{MLP}(\texttt{Flatten}(\mathbf{S}_{t}^{m})),\end{gathered}$ (5)
where $\mathbf{S}_{t}^{m}$ is the intermediate representation of the patch-level similarity, and Flatten indicates the flatten operation. By sending the intermediate representation $\mathbf{S}_{t}^{m}$ to an MLP, the more fine-grained similarities between every two points in the two patches are comprehensively considered, leading to a high-quality patch-level similarity. The similarities between the point query and all sampling points $\mathbf{w}_{t}\in\mathbb{R}^{M}=[w_{t}^{0},w_{t}^{1},\dots,w_{t}^{M-1}]^{\top}$ will work as attention weights to aggregate their corresponding values $\mathbf{V}_{t}\in\mathbb{R}^{M\times D}$, obtaining the spatial querying result $\Delta\mathbf{f}_{t}^{\mathcal{S}}\in\mathbb{R}^{D}$:
[…]
where $\mathbf{V_{t}}$ are the sampled values of $M$ sampling points on $\mathbf{X}_{t}$. After that, the spatial querying results will be utilized as a residual to complete the CCA:
@@
Following TAPTRv2, we also add the attention-based position update (APU) in CCA, to benefit from the attention weights’ resilience to the domain gap without contaminating the point query. This process can be formulated as:
$\begin{gathered}\Delta\mathbf{l}_{t}=\mathbf{O}_{t}^{\top}\otimes\texttt{SoftMax}\left(\texttt{MLP}\left(\mathbf{w}_{t}\right)/\sqrt{D}\right),\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathbf{l}_{t}^{\prime}\Leftarrow\mathbf{l}_{t}^{\prime}+\Delta\mathbf{l}_{t},\end{gathered}$ (8)
@@
3.4 Auto-Triggered Global Matching
In general, point motions are smooth in video. However, in long videos, especially in offline scenarios, scene cut often occurs, leading to sudden large motions. Since the initial position of a point query in the current frame is inherited from the last frame’s prediction, as we have described in Sec.3.1, the sudden large motion will cost many frames for TAPTRv3 to catch up with the target tracking point.
Thus, when a scene cut occurs [^3] We use the off-the-shelf PySceneDetect (Castellano, ) to detect the scene cuts. at the $t$-th frame, we trigger the global matching module to help TAPTRv3 reestablish tracking. Similar to the previous method (Doersch et al., 2023), the global matching module will construct a similarity map $\mathbf{H}_{t}\in\mathbb{R}^{H\times W}$ between the target tracking point and the current frame’s feature map. However, instead of relying on a point-level feature as in previous methods, similar to CCA, we leverage the spatial context features to help improve the accuracy of the similarity map. After that, SoftArgMax is conducted on the similarity map to obtain a final position prediction $\mathbf{l}_{t}\in\mathbb{R}^{2}$ that is differentiable. The prediction will be used to replace the unreliable position prediction of the current frame from the transformer decoder. The replacement helps TAPTRv3 to prevent wrong predictions in the subsequent frames. The process can be formulated as:
@@ 4.1 Datasets and Evaluation
Training Data. For fair comparison, following the previous methods (Doersch et al., 2023; Karaev et al., 2024b; Li et al., 2024c; b), we trained our model on the TAP-Vid-Kubric (Doersch et al., 2022) dataset, which consists of 11,000 synthetic videos generated by Kubric Engine (Greff et al., 2022), each containing 24 frames showing 3D rigid objects falling to the ground and bouncing. Each video has 2,048 points sampled on moving objects and backgrounds to be tracked, and their corresponding trajectories are also generated for training. The points to be tracked in the video are occasionally occluded to allow the model to cope with this situation. During training, the resolution of the videos is resized to $384\times 512$, and we randomly select 800 trajectories in each video for efficient training.
@@
Evaluation Data. We follow previous methods to evaluate TAPTRv3 on the challenging TAP-Vid (Doersch et al., 2022) benchmark, which consists of 3 subsets, TAP-Vid-Kinetics, TAP-Vid-DAVIS, and RGB-Stacking. TAP-Vid-DAVIS comprises 30 real-world videos from the DAVIS 2017 validation set (Perazzi et al., 2016). These videos are relatively short, averaging less than 70 frames, so we include the experiments on this subset in Section A.3 of the appendix. TAP-Vid-Kinetics contains 1,144 YouTube videos from the Kinetics-700-2020 validations set (Carreira & Zisserman, 2017), with camera shakes, complex environments, and even scene cuts. These videos are relatively long, averaging about 250 frames per video. RGB-Stacking is a synthetic dataset that captures the process of a robotic arm grasping solid-colored blocks, with an average duration of about 250 frames. Although it has a relatively smaller domain gap compared to the training set, which is also a synthetic dataset, the objects in this dataset often lack texture, making them difficult to track. In addition, we also evaluated TAPTRv3 on RoboTAP (Vecerik et al., 2024), which comprises 265 real-world videos from robotic manipulation tasks. The video length in this dataset varies significantly, with the longest videos reaching up to 1,300 frames and the average length of about 270 frames per video.
Table 1: Comparison of TAPTRv3 with prior methods. We use to indicate the introduction of additional training data. Specifically, CoTracker3 additionally incorporates 15K real videos. BootsTAPIR and BootsTAPNext-B train on an additional 15M real videos. Anthro-LocoTrackR leverages an extra 1.4K human motion data. TAPTRv3 obtains state-of-the-art performance on most datasets and remains competitive with methods trained on extra internal data. Training data: (Kub24), (Kub48), and (Kub64) refer to Kubric (Greff et al., 2022) with 24, 48, and 64 frames per video, respectively. (PO) PointOdyssey (Zheng et al., 2023), (FT) FlyingThings++ (Mayer et al., 2016). For fair comparison, we do not utilize auto-triggered global matching here.
Training Kinetics RGB-Stacking RoboTAP
Method Data AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA AJ $\delta^{x}_{avg}$ OA
PIPs (Harley et al., 2022) FT 31.7 53.7 72.9 59.1
PIPs++ (Zheng et al., 2023) PO 63.5 58.5 63.0
TAP-Net (Doersch et al., 2022) Kub24 38.5 54.4 80.6 53.5 68.1 86.3 45.1 62.1 82.9
TAPIR (Doersch et al., 2023) Kub24 49.6 64.2 85.0 55.5 69.7 88.0 59.6 73.4 87.0
CoTracker (Karaev et al., 2024b) Kub24 49.6 64.3 83.3 67.4 78.9 85.2 58.6 70.6 87.0
TAPTR (Li et al., 2024c) Kub24 49.0 64.4 85.2 60.8 76.2 87.0 60.1 75.3 86.9
TAPTRv2 (Li et al., 2024b) Kub24 49.7 64.2 85.7 53.4 70.5 81.2 60.9 74.6 87.7
LocoTrack (Cho et al., 2024) Kub24 52.9 66.8 85.3 69.7 83.2 89.5 62.3 76.2 87.1
CoTracker3 (online) (Karaev et al., 2024a) Kub64 54.1 66.6 87.1 71.1 81.9 90.3 60.8 73.7 87.1
Track-On (online) (Aydemir et al., 2025) Kub24 53.9 67.3 87.8 71.4 85.2 91.7 63.5 76.4 89.4
BootsTAPIR (Doersch et al., 2024) Kub24+15M 54.6 68.4 86.5 70.8 83.0 89.9 64.9 80.1 86.3
CoTracker3 (online) (Karaev et al., 2024a) Kub64+15K 55.8 68.5 88.3 71.7 83.6 91.1 66.4 78.8 90.8
Anthro-LocoTrack (Kim et al., 2025) Kub24+1.4K 53.9 68.4 86.4 - - - 64.7 79.2 88.4
BootsTAPNext-B (online) (Zholus et al., 2025) Kub48+15M 57.3 70.6 87.4 - - - - - -
TAPTRv3 (online) Kub24 54.9 67.5 88.2 72.3 84.1 90.8 64.5 77.3 89.7
@@
Evaluation Metrics and Settings. We use the standard metrics proposed in TAP-Vid (Doersch et al., 2022) for evaluation, including three important metrics. Occlusion Accuracy (OA), describes the accuracy of classifying whether a target tracking point is visible or occluded. $\delta_{avg}^{x}$, reflecting the average precision of visible points’ position at thresholds of 1, 2, 4, 8, and 16 pixels. Average Jaccard (AJ), a comprehensive metric, considers both the precision of position and visibility prediction. Since TAPTRv3 is an online tracker, we use the “First query” mode (Doersch et al., 2022) to evaluate the model, which tracks the target tracking point from the first frame when they are visible until the end of the video. This is much more difficult than the “Strided query” mode for offline trackers. Besides, since the resolution of the input video has a large impact on the final performance, we limit the resolution of the input video to $256\times 256$ for a fair comparison with other methods during evaluation.